### v2

# Spatial distribution of organics in the Horsehead nebula: signposts of chemistry driven by atomic carbon

11institutetext: Instituto de Astrofísica, Pontificia Universidad Católica de Chile, Av. Vicuña Mackenna 4860, 7820436 Macul, Santiago, Chile 22institutetext: Millennium Nucleus on Young Exoplanets and their Moons (YEMS) 33institutetext: European Southern Observatory, Alonso de Córdova 3107, Casilla 19, Vitacura, Santiago, Chile 44institutetext: IRAM, 300 rue de la Piscine, 38400 Saint Martin d’Hères, France 55institutetext: LUX, Observatoire de Paris, PSL Research University, CNRS, Sorbonne Universités, 75014 Paris, France 66institutetext: Theoretical Astrophysics, Department of Physics and Astronomy, Uppsala University, Box 516, 75120 Uppsala, Sweden 77institutetext: Instituto de Física Fundamental (CSIC), Calle Serrano 121, 28006, Madrid, Spain 88institutetext: LUX, Observatoire de Paris, PSL Research University, CNRS, Sorbonne Universités, 92190 Meudon, France 99institutetext: Joint ALMA Observatory, Avenida Alonso de Córdova 3107, Vitacura, Santiago, Chile 1010institutetext: Laboratoire d’Astrophysique de Bordeaux, Univ. Bordeaux, CNRS, B18N, Allée Geoffroy Saint-Hilaire, 33615 Pessac, France
Spatial distribution of organics in the Horsehead nebula:
signposts of chemistry driven by atomic carbon
[…]
(Received 25 September 2025 / Accepted 27 January 2026)
Complex organic molecules (COMs) are considered essential precursors to prebiotic species in the interstellar and circumstellar medium. Despite their astrobiological relevance, many aspects of COMs formation remain unclear, particularly the role of ultraviolet (UV) radiation. While COMs were once expected to be efficiently destroyed under UV-irradiated conditions, detections in photodissociation regions (PDRs) have challenged this view. However, the mechanisms by which UV radiation contributes to their formation are still uncertain. Here, we present moderately resolved maps of simple and complex organic molecules at the UV-illuminated edge of the Horsehead nebula, obtained by combining Atacama Large Millimeter/submillimeter Array (ALMA) and IRAM 30m single-dish observations at $\sim$$15^{\prime\prime}$ resolution. For the first time in this PDR environment, we analyze the spatial distribution of species such as C17O, CH2CO, CH3CHO, HNCO, CH3CN, and HC3N. By incorporating previous C17O and C18O single-dish data as well as Plateau de Bure Interferometer (PdBI) maps of H2CO and CH3OH, we derive profiles of gas density, temperature, thermal pressure, and column densities of the organic species as a function of distance from the UV source. Our results show that most organic species—particularly H2CO, CH2CO, CH3CHO, HNCO, and CH3CN—exhibit enhanced column densities at the UV-illuminated edge compared to cloud interiors, possibly indicating efficient dust-grain surface chemistry driven by the diffusion of atomic C and radicals produced via photodissociation of CO and CH3OH, as supported by recent laboratory experiments. The exceptions, HC3N and CH3OH, can be attributed to inefficient formation on dust grains and ineffective non-thermal desorption into the gas phase, respectively. Additionally, contributions from gas-phase hydrocarbon photochemistry—possibly seeded by grain-surface products—cannot be ruled out. Further chemical modeling is needed to confirm the efficiency of these pathways for the studied species, which could have important implications for other cold, UV-irradiated environments such as protoplanetary disks.
Key Words.:
astrochemistry – ISM: clouds – ISM: molecules – ISM: photon-dominated region (PDR)
[…]
1 Introduction
Molecular line observations provide a powerful tool for probing the structure and physical conditions of the molecular gas across a wide range of environments, from high-redshift galaxies (Carilli and Walter 2013) to protoplanetary disks (Öberg et al. 2023). To date, about 330 molecules have been detected in the interstellar medium (ISM) and in circumstellar shells [^1] https://cdms.astro.uni-koeln.de/classic/molecules. While most detections involve simple diatomic or triatomic species, a significant subset consists of organic molecules with six or more atoms, loosely defined as complex organic molecules (COMs; Herbst and van Dishoeck 2009).
The study of COMs and their precursors is critical not only because they trace the physical and chemical evolution across different stages of star and planet formation, but also because they are potential precursors of prebiotic species, the building blocks of terrestrial life (Ceccarelli et al. 2023). Fundamental compounds such as amino acids and nucleotide bases are believed to have an interstellar origin (Garrod 2013; Altwegg et al. 2016; Jiménez-Serra et al. 2020; Rivilla et al. 2019, 2023), though their formation pathways remain uncertain. Therefore, dedicated studies of the formation mechanisms of simpler COMs are essential to clarify their role in producing the ingredients necessary for life as we know it.
The origin of COMs remains an active area of investigation in astrochemistry. Traditionally, COMs have been abundantly detected in hot cores around high-mass protostars (e.g., Bisschop et al. 2007; Feng et al. 2015; Bonfand et al. 2017; van der Walt et al. 2021) and in their low-mass counterparts, hot corinos (e.g., Cazaux et al. 2003; Öberg et al. 2014; Belloche et al. 2020; Yang et al. 2021). Their formation was typically attributed to small organics produced via grain-surface reactions from inherited ices, which are then thermally desorbed and processed further through gas-phase chemistry (Herbst and van Dishoeck 2009). However, the detection of abundant COMs in pre-stellar cores has challenged this view (e.g., Bacmann et al. 2012; Cernicharo et al. 2012; Jiménez-Serra et al. 2016; Scibelli et al. 2024), highlighting the possible importance of cold gas-phase reactions (Shannon et al. 2013), non-diffusive chemistry on grains (Jin and Garrod 2020; Garrod et al. 2022) and non-thermal desorption processes such as chemical desorption (Garrod et al. 2007).
Thus, there is general consensus that the observed abundances of COMs likely arise from an intricate interaction between gas-phase and grain-surface chemistry (Jørgensen et al. 2020; Ceccarelli et al. 2023), though many aspects remain poorly understood. In particular, the role of ultraviolet (UV) radiation is crucial, as it can further process interstellar ices and promote their release into the gas phase through non-thermal photodesorption (Öberg 2016). Moreover, laboratory experiments demonstrate that UV irradiation of ices can even lead to the formation of prebiotic species (de Marcellus et al. 2015; Oba et al. 2019). These processes are especially relevant in environments such as protoplanetary disks, where compelling evidence for in situ ice processing has recently been proposed (e.g., Walsh et al. 2016; Calahan et al. 2023; Ligterink et al. 2024; Yamato et al. 2024; Evans et al. 2025). However, the generally low abundance of COMs and the complex physicochemical structure of these environments (see Öberg et al. 2023, for a recent review) make their study especially challenging.
A promising way to investigate the impact of UV radiation on COM chemistry is to study the edges of molecular clouds illuminated by massive stars, the so-called photodissociation regions (PDRs; Hollenbach and Tielens 1999). In these environments, far-UV (FUV; $6<h\nu<13.6$ eV) photons dominate the chemistry and were initially thought to efficiently photodissociate large molecules. However, more recent studies have challenged this view (see Wolfire et al. 2022, for a recent review). Particularly, COMs have been observed to be abundant in PDRs (e.g., Guzmán et al. 2014; Cuadrado et al. 2017), even when the radiation field exceeds the interstellar average by several orders of magnitude ($G_{0}\sim 10^{2}-10^{4}$, where $G_{0}=1.7$ is the average interstellar radiation field; Draine 1978). Pure gas-phase models cannot account for the observed high COM abundances in these regions (e.g., Gratier et al. 2013; Guzmán et al. 2013; Cuadrado et al. 2017), pointing to a key role for FUV processing of ices. However, the details of this chemistry remain poorly constrained observationally: only a handful of targeted studies exist (e.g., Guzmán et al. 2014; Cuadrado et al. 2017), all relying on single-dish telescopes and thus lacking spatial resolution to determine how COMs are distributed relative to the UV source.
Table 1: Observational parameters for the ALMA-7m+30m maps shown in Figs. 2 and 8.222The projection center of the maps is $\alpha_{2000}=05^{\mathrm{h}}40^{\mathrm{m}}54\hbox{$.\!\!^{\mathrm{s}}$}555$, $\delta_{2000}=-02^{\circ}27^{\prime}48\hbox{$.\!\!^{\prime\prime}$}6$
Species Transition Chan. Width Beam Beam PA Chan. rms Mom. Zero rms
(km s-1) (′′) (mK) (mK km s-1)
C17O J=10J=1-0 $0.20$ $15.69$ $\times$ $8.79$ 73.43-73.43 $55.52$ $27.20$
CH2CO $J_{K_{a},K_{c}}=5_{15}-4_{14}$ $0.20$ $17.08$ $\times$ $9.70$ 77.77-77.77 $34.00$ $16.65$
CH3CHO $J_{K_{a},K_{c}}=5_{14}-4_{13}$ (E) $0.20$ $17.28$ $\times$ $9.85$ 76.92-76.92 $29.26$ $14.33$
$J_{K_{a},K_{c}}=6_{16}-5_{15}$ (E) $0.20$ $15.68$ $\times$ $8.80$ 73.36-73.36 $47.74$ $21.35$
$J_{K_{a},K_{c}}=6_{16}-5_{15}$ (A) $0.20$ $15.68$ $\times$ $8.80$ 73.36-73.36 $47.74$ $21.35$
HNCO $J_{K_{a},K_{c}}=5_{05}-4_{04}$ $0.20$ $15.62$ $\times$ $9.01$ 77.83-77.83 $53.71$ $26.31$
HC3N J=1110J=11-10 $0.20$ $17.08$ $\times$ $9.70$ 77.78-77.78 $35.82$ $17.55$
CH3CN $J_{K}=6_{0}-5_{0}$ (A) $0.20$ $15.52$ $\times$ $8.91$ 78.00-78.00 $63.00$ $35.64$
$J_{K}=6_{1}-5_{1}$ (E) $0.20$ $15.52$ $\times$ $8.91$ 78.00-78.00 $63.00$ $35.64$
$J_{K}=6_{2}-5_{2}$ (E) $0.20$ $15.52$ $\times$ $8.91$ 78.00-78.00 $63.00$ $35.64$
Illuminated by the O9.5V binary system $\sigma$ Orionis (Warren and Hesser 1977) $\sim$$3.5$ pc away from the molecular cloud (Abergel et al. 2003), the PDR located at the edge of the Horsehead nebula is particularly well-suited for a spatially resolved study. Thanks to its proximity ($\sim$$400$ pc, Anthony-Twarog 1982), nearly edge-on geometry, and moderate radiation field ($G_{0}\sim 100$, Habart et al. 2005), the Horsehead PDR is an archetype for the average UV radiation fields in molecular clouds of the Milky Way and normal star-forming galaxies. As a consequence, its gas and dust content have been thoroughly characterized at different wavelengths (Schirmer et al. 2020; Hernández-Vera et al. 2023; Abergel et al. 2024, and references therein). Specifically, the Horsehead WHISPER (Wideband High-resolution IRAM-30m Surveys at two Positions with EMIR Receivers, PI: J. Pety) single-dish line survey provided a deep chemical inventory at two key locations [^3] The PDR is located at $\alpha_{2000}=05^{\mathrm{h}}40^{\mathrm{m}}53\hbox{$.\!\!^{\mathrm{s}}$}936$, $\delta_{2000}=-02^{\circ}28^{\prime}00^{\prime\prime}$ and the dense core is located at $\alpha_{2000}=05^{\mathrm{h}}40^{\mathrm{m}}55\hbox{$.\!\!^{\mathrm{s}}$}61$, $\delta_{2000}=-02^{\circ}27^{\prime}38^{\prime\prime}$.: the HCO emission peak (PDR position, Gerin et al. 2009) and the DCO+ emission peak (dense core position, Pety et al. 2007). Strikingly, this survey yielded the first PDR detections of complex species such as HCOOH, CH2CO, CH3CHO, and CH3CCH (Guzmán et al. 2014). Since then, several smaller molecules have been mapped at higher angular resolution with the Plateau de Bure Interferometer (PdBI; e.g., Guzmán et al. 2013, 2015) and the Atacama Large Millimeter/submillimeter Array (ALMA; e.g., Hernández-Vera et al. 2023), though similar mapping of COMs remains scarce.
In the present study, we investigate the spatial distribution of O-bearing and N-bearing organics at the UV-illuminated edge of the Horsehead nebula, combining ALMA and IRAM 30m data to achieve $\sim$$15^{\prime\prime}$ resolution (see Fig. 1 for an overview). We examine how gas physical conditions vary across the cloud and explore the chemical formation pathways of COMs under moderate FUV irradiation. Section 2 describes the observations and data combination. Section 3 presents the spatial distributions, physical conditions, and column densities of the analyzed species. In Sect. 4, we elaborate on the possible formation pathways of the studied molecules and discuss implications for FUV-driven organic chemistry in analogous environments. The main results and conclusions are summarized in Sect. 5.
2 Observations and data reduction
In this section, we describe the interferometric ALMA deep observations of simple organics and COMs taken at the edge of the Horsehead nebula that were combined with matched IRAM 30m single-dish observations. For the analysis, they were complemented with Plateau de Bure Interferometer (PdBI) legacy data, previously published by Guzmán et al. (2013), and the publicly available WHISPER data.
2.1 ALMA-7m observations
We used the ALMA Atacama Compact Array (hereinafter ALMA-7m) to perform deep observations of the Horsehead nebula at $\sim 15^{\prime\prime}$ between August 1 2017 and October 1 2017, during Cycle 4 (2016.2.00027.S, PI: V.V. Guzmán). There are 22 successful execution blocks (EBs), each lasting for about $1.5$ h, with an on-source time of $0.8$ h. Single pointing observations centered at $\alpha_{2000}=05^{\mathrm{h}}40^{\mathrm{m}}54\hbox{$.\!\!^{\mathrm{s}}$}555$, $\delta_{2000}=-02^{\circ}27^{\prime}48\hbox{$.\!\!^{\prime\prime}$}6$ were carried out in Band 3 with projected baseline lengths between 5 and 49 m for an average of 10 antennas per EB. The primary beamwidth at $100$ GHz is $107.9^{\prime\prime}$ (see Fig. 1, for a reference in physical units, assuming a distance of $400$ pc). The quasars J0510++1800 or J0522-3627 were observed as the bandpass calibrator; J0423-0120, J0522-3627, or Uranus as the absolute flux calibrator; and J0541-0211 as the phase calibrator.
A single spectral setup was defined, covering the 9811398-113 GHz range and divided into nine spectral windows in the frequency division mode (FDM) of the correlator. As listed in Table 2, these windows target the continuum at $3$ mm and the following rotational transitions: carbon monoxide (C17O) J=10J=1-0, ketene (CH2CO) $J_{K_{a},K_{c}}=5_{15}-4_{14}$, acetaldehyde (CH3CHO) $J_{K_{a},K_{c}}=5_{14}-4_{13}$ (only E symmetry) and $6_{16}-5_{15}$ (both E and A symmetries), isocyanic acid (HNCO) $J_{K_{a},K_{c}}=5_{05}-4_{04}$, cyanoacetylene (HC3N) J=1110J=11-10, and methyl cyanide (CH3CN) $J_{K}=6_{0}-5_{0}$ (A), $6_{1}-5_{1}$ (E), and $6_{2}-5_{2}$ (E). Other molecules such as methyl isocyanide (CH3NC) and propynylidyne ($l$-C3H) were also targeted. The former was not detected, whereas the latter was, but its analysis lies beyond the scope of this work.
Due to the small aggregate bandwidth, the bandwidth-switching technique was adopted in order to improve the signal-to-noise ratios of the flux and gain calibration. Another spectral setup was defined for the calibrator scans. The calibration setup includes four wide-band spectral windows in the time division mode (TDM), centering at roughly $98.32$, $100.28$, $110.32$, and $112.32$ GHz with the channel spacing of $15.625$ MHz and the bandwidth of $2.0$ GHz. In each EB, the bandpass calibrator [^4] This is now referred to as the differential gain calibrator and is observed with the DIFFGAIN scan intent in recent cycles. was observed in both FDM (narrow-band) and TDM (wide-band) setups to determine the phase differences between the FDM and TDM windows (see ALMA Technical Handbook for details, e.g. Cortes et al. 2025), while the flux and phase calibrators were observed with the TDM setup only.
The data were initially calibrated manually at the Joint ALMA Observatory (JAO) using the Common Astronomy Software Application (CASA) [^5] https://casa.nrao.edu/ version 5.1.1 (McMullin et al. 2007; CASA Team et al. 2022) with the calibration scripts created by the ALMA Script Generator (Petry et al. 2014). We have modified the scripts and re-calibrated the data with CASA version 5.4.0 in order to optimize the spectral window mapping schemes for the system temperature ($T_{\rm sys}$) and gain (amplitude and phase) calibrations. The $T_{\rm sys}$ of the calibrators and science target data were derived from the measurement scans of the same source and spectral window. The gain solutions for the narrow-band windows of the target were derived from the wide-band windows of the phase calibrator in the same baseband (see Table 2). In addition, antenna CM06 was manually flagged due to unreliable $T_{\rm sys}$ values. After flagging, the on-source visibilities amount to about 14 hours of integration times with 10 antennas.
Figure 2: Zeroth-moment maps gallery for different molecular species moderately resolved at the Horsehead edge. Maps have been rotated $14^{\circ}$ counterclockwise to bring the illuminating star direction in the horizontal direction. Contours are $3,5,10,15,20,25,30]\times\sigma$, where $\sigma$ is the zeroth-moment rms listed in Table 1. The red vertical line represents the horizontal zero, delineating the PDR edge (Pety et al. 2005), whereas the cyan and green crosses show the dense core (Pety et al. 2007) and PDR (Gerin et al. 2009) positions, respectively. The beam size and a scale bar indicating $4000$ au are shown in the bottom left and bottom right corner, respectively, of each panel. Transitions corresponding to the same molecular species are shown using the same color-scale range.
2.2 IRAM 30m observations
As part of project 108-17, we used the EMIR receivers of the IRAM 30m telescope to map during 47 hours a region of $180^{\prime\prime}\times 180^{\prime\prime}$ centered on ALMA-7m pointing and rotated by $14^{\circ}$ counter-clockwise compared to the Equatorial $2000$ coordinate frame. We used the narrow mode of the Fourier Transform Spectrometer that delivers four times $1.8$ GHz of continuous bandwidths (2 of them in each sideband) at a channel spacing of $50$ kHz. We interleaved two different receiver tunings to cover continuously from the lower sideband from $93.65$ to $100.55$ GHz, and the upper sideband from $109.65$ to $116.65$ GHz. We used the position-switch, on-the-fly observing mode, covering each field with back-and-forth scans along either the right ascension or declination axes. We slewed at a speed of $8.5^{\prime\prime}$ s-1 and we dumped data to disk every $0.5$ s, yielding at least 5 integrations per beam in the scanning direction. The scan legs were separated by $8.5^{\prime\prime}$, yielding Nyquist sampling transverse to the scan direction at $2.6$ mm. We calibrated the system every $15$ minutes.
The calibration of the data was done inside the GILDAS/MRTCAL software and the remaining of the processing happened in the GILDAS/CLASS software. Around the frequency of each line of interest, we extracted a frequency range of $100$ MHz. We baselined the spectra by fitting a polynomial of order 1 after excluding the [$8.5$, $12.5$ km s-1] velocity range. We converted the spectra intensity in main beam temperature using the standard Ruze formula for the IRAM 30m telescope. Finally we resampled the velocity axis to match the ALMA-7m one, and we gridded the data in a cube of $4^{\prime\prime}\times 4^{\prime\prime}$ pixels.
Figure 3: Left column: Integrated line intensity profiles along the direction of the exciting star extracted from the maps shown in Fig. 2 and the convolved maps obtained from Guzmán et al. (2013). For molecules with more than one transition, the brightest is displayed. The profiles were extracted at the $\delta y$ position of the PDR (top panel) and dense core (bottom panel), and averaged over $10^{\prime\prime}$ in the $\delta y$ direction. The dotted vertical lines represent the $\delta x$ position of the PDR and dense core, and the vertical red line represents the PDR edge. The average beam size is represented by the horizontal gray bar. The colored area of each profile displays the $\pm\sigma$ significance levels, taking into account the standard deviation of the average and the zeroth-moment rms. Right column: Regions used to extract the profiles, indicated by white rectangles overlaid on the HNCO zeroth-moment map. The two crosses, the red vertical line, and the contours are the same as in Fig. 2.
2.3 Interferometric and single-dish data combination
We combined the IRAM 30m and ALMA-7m to produce a clean image containing all spatial information from 0 to $50$ m, i.e., equivalent to a total power telescope of 5050m-diameter. The IRAM 30m map is deconvolved from the $30$ m beam in the Fourier plane before multiplication by the ALMA-7m primary beam in the image plane. After a last Fourier transform, pseudo-visibilities were sampled between 0 and $7$ m. These visibilities were then merged with the ALMA-7m interferometric observations (see Rodriguez-Fernandez et al. 2008, for details). Standard imaging, CLEAN deconvolution, and primary beam correction steps were then applied. The resulting data cube was then scaled from Jy beam-1 to $T_{\mathrm{mb}}$ temperature scale using the synthesized beam size. All these processes happened in the GILDAS/MAPPING software.
The observational parameters of the final combined (hereinafter ALMA-7m+30m) images are summarized in Table 1. The channel rms is calculated using the central 50% of the pixels in the first and last five (line-free) channels of the data cubes. To compare the spatial distribution of the different molecular tracers, we collapse the line emission into velocity-integrated intensity, or zeroth-moment, maps using the CASA task immoments. The zeroth-moment rms is estimated as $\sigma_{\mathrm{rms}}\approx\sigma_{\mathrm{chan}}\sqrt{N_{\mathrm{chan}}}\Delta v$, where $\sigma_{\mathrm{chan}}$ is the channel rms, $N_{\mathrm{chan}}$ is the number of channels used to compute the zeroth-moment, and $\Delta v$ is the channel spacing in velocity.
2.4 Legacy data
The analysis of the ALMA-7m+30m data was complemented with PdBI observations of formaldehyde (H2CO) $J_{K_{a},K_{c}}=2_{02}-1_{01}$ and methanol (CH3OH) $J_{K}=3_{-1}-2_{-1}$ (E symmetry) rotational transitions imaged at $\theta_{\mathrm{beam}}\sim 6^{\prime\prime}$. All the observational details of these legacy data can be found in Guzmán et al. (2013). In order to make the two data sets comparable, the PdBI data was rotated by 14, and the angular resolution was downgraded to $\theta_{\mathrm{beam}}\sim 15^{\prime\prime}$ to match the ALMA-7m+30m data using the CASA tools rotate and convolve2d, respectively. Afterward, the spatial axes were regridded using the CASA task imregrid to match the number of pixels per beam. Throughout this procedure, the CH3CN image was used as a template. Thus, for all subsequent analysis, the angular resolution of the H2CO and CH3OH images is exactly the same as for CH3CN (see Table 1).
Additionally, we extracted from the WHISPER line survey [^6] https://iram.fr/~pety/horsehead/Horsehead_Nebula/Data.html the spectra of the J=10J=1-0 ($\theta_{\mathrm{beam}}\sim 22^{\prime\prime}$) and 212-1 ($\theta_{\mathrm{beam}}\sim 11^{\prime\prime}$) rotational transitions for the C17O and C18O isotopologues at the PDR and dense core positions, to estimate the $\mathrm{C}^{17}\mathrm{O}/\mathrm{C}^{18}\mathrm{O}$ ratio in the Horsehead nebula and find the excitation temperature of C17O (see Sect. 3.2.1 and Appendix B.1). The observational details of WHISPER can be found in any of the survey papers (e.g., Guzmán et al. 2012; Pety et al. 2012).
3 Results
This section presents the main findings from our combined ALMA-7m+30m observations. We describe the spatial distribution of simple and complex organics in the Horsehead nebula, based on integrated intensity maps and their corresponding horizontal profiles at key vertical positions. Next, we detail the radiative transfer modeling used to constrain the physical conditions and derive column density profiles for each organic species. Finally, we compare our results with previous studies in the same source and discuss the differences and similarities among the species analyzed.
3.1 Spatial distribution of molecules
Figure 2 shows the gallery of zeroth-moment maps for all the molecular species listed in Table 1, except for C17O that is shown in Fig. 8. Following the usual conventions of previous work, the maps were rotated $14^{\circ}$ counterclockwise to bring the exciting star direction in the horizontal $\delta x$ direction. By doing this, a new rotated coordinate system is defined where the horizontal zero ($\delta x=0$) corresponds to the empirical PDR edge (red line, Fig. 2) usually defined at the sharp boundary of the molecular cloud traced by the H2 ro-vibrational line emission (Pety et al. 2005). Using the rotated coordinates, the PDR (green cross, Fig. 2) and dense core (cyan cross, Fig. 2) are located at ($\delta x\approx 15\hbox{$.\!\!^{\prime\prime}$}1,\>\delta y\approx 1\hbox{$.\!\!^{\prime\prime}$}2$) and ($\delta x\approx 44\hbox{$.\!\!^{\prime\prime}$}7,\>\delta y\approx 16\hbox{$.\!\!^{\prime\prime}$}5$), respectively.
The visual inspection of Fig. 2 reveals a distinct trend: most molecules, except for HC3N, show a filament structure whose emission peak is at the PDR position. However, it is also clear that O-bearing and N-bearing molecules apparently have different spatial distributions, which probably implies disparities in their chemistry (see Sect. 4, for a detailed discussion). In addition to exhibiting a filament at the PDR, O-bearing species, such as CH2CO and CH3CHO, have an extended component towards the position of the dense core, following the shape of the “Horsehead pillar” typically shaped by large-scale structure maps of CO isotopologues (Roueff et al. 2024; Ségal et al. 2024, see also Fig. 8). Instead, N-bearing species, such as CH3CN and HC3N, are notably different. CH3CN lacks significant extended emission at the dense core and the PDR filament is thinner than the O-bearing ones. In contrast, HC3N reveals a more clumpy morphology concentrated close to the dense core. In between, we have HNCO, exhibiting characteristics of both types of species: a sharply defined filament across the PDR along with some extended emission near the dense core, consistent with its dual N and O composition.
Interestingly, both the filament centered at the PDR and the clumps close to the dense core have been previously seen by higher angular resolution data of other molecules. Guzmán et al. (2013) showed that H2CO emission presents a similar morphology, and suggested it is a consequence of dust-grain versus gas-phase chemistry at the PDR and dense core, respectively. Consistently, Hernández-Vera et al. (2023) concluded that HCO+ exhibits a similar type of distribution since it traces the dense UV-shielded gas but also the more diffuse edge of the molecular cloud exposed to stellar UV radiation. Thus, the different distributions of CH3CN and HC3N seen in the maps could be direct evidence of different formation pathways, as previously suggested by Gratier et al. (2013), or variations on the gas-phase C/O ratio, as mentioned by Le Gal et al. (2019). Likewise, similarities between CH2CO and CH3CHO might be associated with common formation routes, as proposed by Guzmán et al. (2014). More details on the possible formation pathways are discussed in Sect. 4.
Similar to previous works (e.g., Guzmán et al. 2015; Hernández-Vera et al. 2023), we extracted horizontal cuts from the maps to quantify the distribution of the line emission as a function of $\delta x$, particularly at the $\delta y$ position of the PDR and dense core. In this case, we averaged the integrated line intensity over $10^{\prime\prime}$ in the $\delta y$ direction since it roughly matches the beam of the observations in that direction (see right panels, Fig. 3). Hereinafter, we decided to include the convolved maps of H2CO and CH3OH in the analysis to compare the emission of these smaller species with the larger COMs presented here. Moreover, they can be essential precursors of more complex species (Öberg et al. 2009; López-Sepulcre et al. 2024).
Figure 3 shows the integrated intensity profiles extracted at $\delta y\approx 1\hbox{$.\!\!^{\prime\prime}$}2$ (PDR) and $\delta y\approx 16\hbox{$.\!\!^{\prime\prime}$}5$ (dense core). The uncertainties of each profile incorporate the standard deviation of the average along the $\delta y$ direction and the zeroth-moment rms from Table 1, added in quadrature. We also accounted for an absolute flux calibration uncertainty of approximately 5%. For H2CO and CH3OH, the zeroth-moment rms is estimated from the observational parameters reported by Guzmán et al. (2013). It is now more evident that most molecules shown in Fig. 2 have their emission peak at the PDR filament and that the N-bearing molecules CH3CN and HC3N have different spatial distributions. Hence, to further explore the above, we used the integrated intensity profiles to determine the gas physical conditions and the column densities profiles for each molecule.
3.2 Radiative transfer modeling
We have performed radiative transfer calculations to model the line intensities shown in Fig. 3. Typically, the usual way of doing this is through a rotational diagram analysis (Goldsmith and Langer 1999), which assumes optically thin emission and level populations following a Boltzmann distribution characterized by a single excitation temperature ($T_{\mathrm{ex}}$) and the total column density of the molecule ($N$). Under local thermodynamic equilibrium (LTE), $T_{\mathrm{ex}}$ provides a good approximation of the gas kinetic temperature ($T_{\mathrm{kin}}$). However, previous studies have shown that complex molecules in the Horsehead edge are, in general, sub-thermally excited ($T_{\mathrm{ex}}<T_{\mathrm{kin}}$; see Guzmán et al. 2014, and references therein). Therefore, to obtain more accurate results, it is always desirable to take a non-LTE approach, in which both collisional and radiative de-excitation must be considered.
We used the non-LTE radiative transfer code RADEX (van der Tak et al. 2007). More specifically, we ran our analyses using the Python-wrapped version SpectralRadex [^7] https://github.com/uclchem/SpectralRadex (Holdship et al. 2021). Taking this approach, the modeled emission of a specific molecule (in units of K km s-1) will depend on three free parameters: the gas density of the collisional partners, typically H2 ($n_{\mathrm{H_{2}}}$) [^8] We neglect the possible influence of H, He, and e- as collisional partners., the gas temperature ($T_{\mathrm{kin}}$), and the total column density of the molecule ($N$). Other parameters, such as the background temperature ($T_{\mathrm{bg}}$) and the linewidth ($\Delta v$), can also be varied. For simplicity, and because they are not expected to vary significantly across the analyzed region, we kept them fixed at $T_{\mathrm{bg}}=2.73$ K (CMB temperature) and $\Delta v=0.6$ km s-1, as supported by previous studies (Gratier et al. 2013; Guzmán et al. 2014).
We fitted the modeled emission to the observed integrated intensities by constructing a likelihood function, $\mathcal{L}(\mathrm{data},\mathrm{free\>parameters})$, and performing a $\chi^{2}$ minimization. To retrieve the posterior probability distributions of the free model parameters, we employed the affine-invariant Markov Chain Monte Carlo (MCMC) algorithm implemented in the emcee package (Foreman-Mackey et al. 2013). The number of walkers was set to four times the number of free model parameters, with 2000 burn-in steps and an additional 3000 steps used to sample the posterior distributions. The observational uncertainties used in the fit correspond to those shown in the profiles in Fig. 3. The best-fit values and their associated uncertainties were derived from the 16th, 50th, and 84th percentiles of the posterior distributions.
The number of free model parameters in the fit depended on the number of observed transitions available for each molecule. For CH3CN, which has multiple detected rotational lines, we allowed $T_{\mathrm{kin}}$ and $N$ to vary freely. For the other molecules, which had only one or two detected lines, we fixed some of the physical parameters (typically $n_{\mathrm{H_{2}}}$ and/or $T_{\mathrm{kin}}$) based on the results from CH3CN or literature values, and then constrained the column density $N$ through the fitting. To avoid oversampling given the $\sim$$15^{\prime\prime}$ beam size, the full analysis was carried out at intervals of $\sim$$7.5^{\prime\prime}$, ensuring that each measurement is separated by approximately half a beam.
3.2.1 C17O-to-H2 conversion factor and $N(\mathrm{C^{18}O})/N(\mathrm{C^{17}O})$ ratio
One key parameter for the non-LTE radiative transfer analysis is $n_{\mathrm{H_{2}}}$. As the second most abundant molecule in the universe, and given the difficulties of directly observing H2, CO has been historically used to trace the amount of cold molecular gas (Bolatto et al. 2013). Since CO emission is typically optically thick, observations from less abundant isotopologues are more desirable. Nevertheless, sometimes even 13CO emission is optically thick, and thus the use of rarer isotopologues, such as C17O, is needed. Since we have prior information about $n_{\mathrm{H_{2}}}$ at the PDR and dense core, we first fitted the C17O emission at these two specific positions to calibrate the C17O-to-H2 conversion, and then extended it to the other $\delta x$ positions to derive an $n_{\mathrm{H_{2}}}$ profile.
Although the collisional rates of C17O with H2 are available (Yang et al. 2010), we decided to model the emission using the LTE approach. This is a reasonable assumption, given that the critical density of the C17O J=10J=1-0 transition is generally $n_{\mathrm{crit}}<n_{\mathrm{H_{2}}}$ (see Table 3) for the typical gas densities ($n_{\mathrm{H_{2}}}$ $\sim 10^{4}-10^{5}$ cm-3) and temperatures ($T_{\mathrm{kin}}$ $\lesssim 100$ K) observed in the inner layers of the Horsehead nebula (Habart et al. 2005; Pety et al. 2005). Furthermore, we modeled the spectral lines directly, rather than the integrated intensity, to properly account for the hyperfine structure of the emission. Following the procedure described in Hernández-Vera et al. (2024), the line intensity ($I_{\nu}$) in LTE can be expressed in terms of $N$, $T_{\mathrm{ex}}$, and $\Delta v$. However, since we only had one C17O transition observed with ALMA-7m+30m, it was not possible to simultaneously constrain all three parameters.
Therefore, taking advantage of its wide spectral coverage, we used WHISPER data to simultaneously fit the C17O 101-0 and 212-1 transitions. We also fitted the same transitions for C18O to determine the C18O/C17O ratio and verify the consistency of our C17O conversion factor with other CO isotopologues. All relevant spectroscopic constants used for the hyperfine structure fitting are provided in Table 3, and the best-fit models are shown in Fig. 7. Each hyperfine component was modeled as an independent Gaussian with fixed velocity separations, given that they are partially blended at the $0.2$ km s-1 resolution. Their relative intensities were set by the ratios of the upper-state degeneracies, $g_{u}$, and the Einstein coefficients, $A_{ul}$. From these results, we derive an average isotopologue ratio of $N(\mathrm{C^{18}O})/N(\mathrm{C^{17}O})=3.6\pm 0.5$, in good agreement with the local ISM 18O/17O element ratio (3.2, Wilson and Rood 1994).
Interestingly, a $T_{\mathrm{ex}}$ value of <20<20 K for C17O and C18O appears to be appropriate for both the PDR and dense core positions. However, given the low $n_{\mathrm{crit}}$ of their rotational transitions, LTE conditions would imply $T_{\mathrm{ex}}\approx T_{\mathrm{kin}}=60$ K for the PDR (Gerin et al. 2009). Since the emission is barely resolved, the beam likely includes contributions from the colder, denser cloud interior rather than solely from the PDR. To test this, we repeated the analysis with SpectralRadex, fixing $n_{\mathrm{H_{2}}}$ to values from previous work (Gerin et al. 2009), and obtained similarly low $T_{\mathrm{kin}}$ and $T_{\mathrm{ex}}$ values ($\sim 13-20$ K), revealing beam pickup from colder material within the cloud that is likely in LTE.
Hence, with the excitation temperature fixed at $T_{\mathrm{ex}}=15$ K, we repeated the fitting procedure but now for the ALMA-7m+30m data, leaving $N$ and $\Delta v$ as free parameters. This approach allows us to determine NN(C17O)—needed to estimate $n_{\mathrm{H_{2}}}$—while $\Delta v$ is constrained by fitting the hyperfine components of the observed transition. The best-fit models for the PDR and dense core positions are shown in Fig. 8. Defining the C17O-to-H2 conversion factor as $X{{}_{\mathrm{C^{17}O}}}=N(\mathrm{H_{2}})/N(\mathrm{C^{17}O})$, we combined the $N(\mathrm{H_{2}})$ estimates for the PDR and dense core from Guzmán et al. (2014) with our $N(\mathrm{C^{17}O})$ measurements to derive empirical conversion factors for the Horsehead nebula. The resulting values are $X_{\mathrm{C^{17}O}}=(4.0\pm 0.2)\times 10^{6}$ for the PDR and $X_{\mathrm{C^{17}O}}=(5.9\pm 0.3)\times 10^{6}$ for the dense core.
Being a less abundant isotopologue, C17O is not commonly used as a tracer of H2 column density. As a consistency check, we used our newly derived $N(\mathrm{C^{18}O})/N(\mathrm{C^{17}O})$ ratio to compare our $X_{\mathrm{C^{17}O}}$ conversion factor with values reported in the literature. Assuming $N(^{13}\mathrm{CO})/N(\mathrm{C^{18}O})\sim 8$ (Roueff et al. 2024; Ségal et al. 2024), $N(\mathrm{C^{18}O})/N(\mathrm{C^{17}O})\sim 3.6$ (this work), and $X_{\mathrm{C^{17}O}}=5.0\times 10^{6}$ (average value), we determined a $^{13}\mathrm{CO}-$to-H2 conversion factor of $X_{\mathrm{{}^{13}CO}}\sim 1.7\times 10^{5}$, which is somewhat lower than previous results in the Orion B molecular cloud ($4.0\times 10^{5}$, Roueff et al. 2021) and other typical ISM estimates ($3.8\times 10^{5}$, Bolatto et al. 2013). Since we are observing a PDR environment, this discrepancy is likely due to more efficient selective photodissociation of C17O and C18O compared to the more abundant isotopologue 13CO, as both C17O and C18O become abundant only deeper into the cloud, well behind the H/H2 transition.
3.2.2 Gas density based on C17O observations
Once the $X_{\mathrm{C^{17}O}}$ conversion factor was established, we extended the same fitting procedure of C17O ALMA-7m+30m data to all other $\delta x$ positions at the $\delta y$ locations of the PDR and dense core, using the range shown in Fig. 3. This yielded $N(\mathrm{C^{17}O})$ profiles that we converted into $n_{\mathrm{H_{2}}}$ profiles. The advantage of this method is that, by using ALMA-7m+30m data, we obtain empirical gas density profiles at the same angular resolution as the other molecular line observations, thus enabling more accurate column density determinations for the organic species analyzed in this work.
Assuming a typical cloud depth along the line of sight of $\ell_{\mathrm{depth}}\simeq 0.1$ pc (Habart et al. 2005; Lis et al. 2025), and accounting for the fact that the empirical $X_{\mathrm{C^{17}O}}$ value varies between the PDR and dense core, we define the conversion from NN(C17O) to $n_{\mathrm{H_{2}}}$ as follows:
$n_{\mathrm{H_{2}}}(\delta x)=\frac{X_{\mathrm{C^{17}O}}(\delta x)N(\mathrm{C^{17}O})(\delta x)}{\ell_{\mathrm{depth}}},$ (1)
where $X_{\mathrm{C^{17}O}}(\delta x)$ is obtained by linear interpolation and extrapolation of the empirical conversion factors determined at the positions of the PDR and dense core. This approach preserves consistency with previous $n_{\mathrm{H_{2}}}$ determinations at these positions, while providing a smoothly varying profile between them.
The top panel of Fig. 4 shows the resulting $n_{\mathrm{H_{2}}}$ profiles for the PDR and dense core vertical positions, along with the profile modeled by Habart et al. (2005) convolved with a $15^{\prime\prime}$ FWHM Gaussian to facilitate comparison with the ALMA-7m+30m data. Both profiles confirm the presence of a steep density gradient, with typical values of $n_{\mathrm{H_{2}}}\sim 10^{4}-10^{5}$ cm-3. Although lower densities in the surface layers of the cloud are also plausible ($n_{\mathrm{H_{2}}}\lesssim 10^{3}$ cm-3, Hernández-Vera et al. 2023), non-LTE effects and beam pickup from colder and denser regions could still bias the inferred C17O column densities upward, and thus the $n_{\mathrm{H_{2}}}$ values derived at the edge of the cloud.
Figure 4: Empirical gas density (top panel), kinetic temperature (middle panel), and thermal pressure (bottom panel) profiles determined from C17O and CH3CN observations of the Horsehead edge, extracted at the $\delta y$ position of the PDR (blue circles) and dense core (orange circles). The beam sizes of the respective molecular tracers from which the values were derived are represented by the horizontal gray bar in the top left corner of each panel. The red vertical line and the dotted vertical lines are the same as in the left panels of Fig. 3. For comparison, the gas density profile modeled by Habart et al. (2005) convolved with a Gaussian of 15′′ FWHM (green line) and the thermal pressure range from Hernández-Vera et al. (2023) (purple-shaded region) are also shown.
3.2.3 Gas temperature based on CH3CN observations
CH3CN is often considered as a “good thermometer” due to its properties of symmetric top molecule (e.g., Guesten et al. 1985). Considering that there cannot be radiative transitions between different $K$-ladders, the energy levels with different $K$ values are populated by collisions. Therefore, the relative intensities between transitions with the same $J$ but different $K$ values are sensitive to $T_{\mathrm{kin}}$.
As described at the beginning of Sect. 3.2, we used SpectralRadex to model the integrated line intensity of the three CH3CN J=65J=6-5 $(K=0,1,\mathrm{and}\>2)$ transitions simultaneously. In this case, we decided to employ the non-LTE approach since Gratier et al. (2013) demonstrated that CH3CN emission at the Horsehead PDR is sub-thermally excited. The collisional rates were taken from Ben Khalifa et al. (2023), and we assumed a ratio of $\mathrm{E/A}=1$ between the two non-interacting torsional substates. Using the $n_{\mathrm{H_{2}}}$ values derived from C17O as input, the modeled CH3CN intensities only depend on $T_{\mathrm{kin}}$ and $N(\mathrm{CH_{3}CN})$ as free parameters. The best-fit integrated intensity models are shown in Fig. 9. The resulting $T_{\mathrm{kin}}$ profiles for the PDR and dense core vertical positions are shown in the middle panel of Fig. 4, whereas the $N(\mathrm{CH_{3}CN})$ profiles are shown in Fig. 5 and Fig. 10.
We have obtained temperature values of $T_{\mathrm{kin}}=58^{+13}_{-10}$ K for the PDR position ($\delta x\approx 15\hbox{$.\!\!^{\prime\prime}$}1,\>\delta y\approx 1\hbox{$.\!\!^{\prime\prime}$}2$) and $T_{\mathrm{kin}}=28^{+11}_{-10}$ K for the dense core position ($\delta x\approx 44\hbox{$.\!\!^{\prime\prime}$}7,\>\delta y\approx 16\hbox{$.\!\!^{\prime\prime}$}5$), which is in good agreement with previous estimates using CH3CN (Gratier et al. 2013) and other molecular tracers (Pety et al. 2007; Gerin et al. 2009; Guzmán et al. 2013). Regarding the other profile positions, we have empirically demonstrated the presence of a temperature gradient, as earlier pointed out by observations (Pety et al. 2005) and models (Goicoechea et al. 2009a; Le Gal et al. 2017, 2019). In this case, the temperature derived from CH3CN at the PDR is significantly higher than that estimated from C17O using WHISPER observations. Although the beam sizes are not very different in the $\delta x$ direction, the spatial distribution of C17O is much more extended than that of the CH3CN filament, which shows almost no counterpart in the dense core and likely minimizes beam-pickup effects from colder interior regions.
However, we did not recover the high temperatures ($T_{\mathrm{kin}}\gtrsim 100$ K) expected at the very edge of the cloud (Zannese et al. 2025a), in part because the collisional rates used were tabulated only up to $100$ K. Still, the posterior probability distributions showed no indication of a preference for higher temperatures. This discrepancy likely stems from the limited angular resolution and sensitivity of the observations, which may prevent detection of the outermost regions of the cloud, characterized by lower column densities and, consequently, lower surface brightness.
3.2.4 Gas thermal pressure
Having derived empirical profiles for the gas density and kinetic temperature, we computed the corresponding thermal pressure, defined as $P_{\mathrm{th}}=n_{\mathrm{H}}T_{\mathrm{kin}}$, where $n_{\mathrm{H}}=n(\mathrm{H})+2n(\mathrm{H_{2}})\sim 2n_{\mathrm{H_{2}}}$ corresponds to the total hydrogen nuclei density. The resulting $P_{\mathrm{th}}$ profiles for the vertical positions of the PDR and dense core are shown in the bottom panel of Fig. 4, together with the range of values determined from high angular resolution observations by Hernández-Vera et al. (2023) [^9] Note that in Hernández-Vera et al. (2023) the factor of 2 in the definition of $n_{\mathrm{H}}\sim 2n_{\mathrm{H_{2}}}$ is missing. As a result, and for consistency with previous work, the range of values used for comparison here is $P_{\mathrm{th}}=(4.6$$8.0)\times 10^{6}$ K cm-3..
These results corroborate the presence of high thermal pressures at the Horsehead edge, exceeding those measured in the adjacent IC 434 H II region (Bally et al. 2018), in agreement with a photoevaporative compression scenario. Moreover, to first order, the Horsehead PDR appears to maintain an isobaric structure: although some variations are present, most values remain, within uncertainties, consistent with the range reported by Hernández-Vera et al. (2023). These findings add compelling new observational support to previous evidence that PDRs can often be represented as isobaric regimes with large thermal pressures regulated by the intensity of the incident UV radiation field (Goicoechea et al. 2016; Joblin et al. 2018; Wu et al. 2018; Bron et al. 2018; Maillard et al. 2021; Hernández-Vera et al. 2023; Clark et al. 2025).
Figure 5: Column density profiles derived at the $\delta y$ position of the PDR from the radiative transfer modeling of the O-bearing (left panel) and N-bearing (right panel) molecules analyzed in this work. For each color, the solid line depicts the best-fit values, whose uncertainties are represented by the shaded areas of the same color. The red vertical line, the dotted vertical lines, and the horizontal gray bar are the same as in the left panels of Fig. 3. The same profiles but at the $\delta y$ position of the dense core are shown in Fig. 10.
3.2.5 Column density profiles
The integrated line emission of the remaining molecules was modeled to obtain column density profiles and further investigate the formation pathways of the different chemical species. Although it was mentioned that complex molecules in the Horsehead nebula are sub-thermally excited (i.e., $n_{\mathrm{crit}}<n_{\mathrm{H_{2}}}$, see Table 3), we cannot use the RADEX non-LTE approach for all the species presented in this work since not all have collisional rates available in the literature.
Thus, for those species with available rates, such as H2CO (Wiesenfeld and Faure 2013), CH3OH (Rabli and Flower 2010), HNCO (Sahnoun et al. 2018), and HC3N (Faure et al. 2016), we performed a similar procedure as CH3CN but using the $n_{\mathrm{H_{2}}}$ and $T_{\mathrm{kin}}$ profiles derived in the previous sections as input. Instead, for CH2CO and CH3CHO, we used the rotational diagram approach by assuming a Boltzmann distribution of the rotational levels characterized by a single $T_{\mathrm{ex}}$ lower than $T_{\mathrm{kin}}$, consistent with a sub-thermal excitation regime. Under this assumption, the column density in the upper state of a given transition ($N_{u}$) is given by:
$N_{u}=\frac{N}{Q(T_{\mathrm{ex}})}g_{u}e^{-E_{u}/kT_{\mathrm{ex}}},$ (2)
where $N$ is the total column density, $Q$ is the partition function of the molecule taken from the JPL database (Pickett et al. 1998), and $E_{u}$ with $g_{u}$ are the energy and the degeneracy associated to the upper-state level of the corresponding transition, respectively (see Table 3). On the other hand, assuming optically thin emission, $N_{u}$ is related to the integrated line intensity ($I=\int I_{\nu}d\nu$, in K km s-1) through the equation:
$I=\frac{hc^{3}A_{ul}}{8\pi k\nu^{2}}N_{u},$ (3)
where $\nu$ is the line frequency and $A_{ul}$ is the Einstein coefficient (see Table 3). By putting Eq. (2) into Eq. (3), the integrated line intensity can therefore be written as a function of $N$ and $T_{\mathrm{ex}}$. Similar to the C17O case, we fixed the excitation temperature at $T_{\mathrm{ex}}=15$ K for CH2CO and $T_{\mathrm{ex}}=8$ K for CH3CHO, based on previous work (Guzmán et al. 2014).
Hence, either using the non-LTE approach implemented by SpectralRadex or the rotational diagram approach through Eqs. (2) and (3), the only free parameter to fit is $N$. The fitting strategy was essentially the same as for CH3CN, which is illustrated in Appendix B.3. It is worth mentioning that, similar to the case of CH3CN, we assumed $\mathrm{E/A}=1$ for CH3OH and CH3CHO and an ortho-to-para ratio of $\mathrm{OPR}=3$ for H2CO and CH2CO since we do not have any information about these symmetry ratios and nuclear spin forms in positions other than the PDR and the dense core. In these cases, the total column density corresponds to the sum of the different symmetries (E+A) or nuclear spin forms (ortho+para).
Figure 5 shows the derived column density profiles for the PDR vertical position, while the corresponding profiles for the dense core vertical position are presented in Fig. 10. Since the profiles at both positions are very similar, we will refer only to Fig. 5 in the subsequent analysis, with the understanding that the same conclusions apply to the dense core vertical position. The associated uncertainties correspond to the statistical propagation of observational errors, without including systematic uncertainties on the fixed parameters, as this would require exploring a multidimensional parameter space in density and temperature, which is beyond the scope of this work. We checked the consistency of our results by comparing them with previous efforts. Since this is the first time that the column density of these molecules is spatially imaged, the only benchmarks that we have are the values derived at the PDR and dense core positions from prior single-dish observations, except for HNCO for which there are no results reported in the literature.
For the molecules observed with the PdBI, our results are consistent within the uncertainties with the values determined by Guzmán et al. (2013) using the IRAM 30m spectra only, confirming that the emission of H2CO and CH3OH is only marginally diluted. Instead, for the combined ALMA-7m+30m maps, most of our column densities are higher than previous results from the WHISPER line survey (Gratier et al. 2013; Guzmán et al. 2014), specifically by a factor of $\sim 3-4$ in the dense core and a factor of $\sim 3-6$ in the PDR. Beam dilution effects can explain the above since, from the maps shown in Fig. 2, we can see that we begin to resolve filamentary and clumpy structures with angular widths on the order of $\theta_{\mathrm{source}}\sim 10-15^{\prime\prime}$ that probably were diluted by the large beam size reached by WHISPER ($\theta_{\mathrm{beam}}\gtrsim 27^{\prime\prime}$, Guzmán et al. 2014).
For CH3CN, at the PDR position we found a column density $\sim 0.5$ times lower than that reported by Gratier et al. (2013) using WHISPER data. However, Gratier et al. (2013) already accounted for beam dilution in their radiative transfer modeling, which explains why we find a similar column density with the new higher-angular resolution observations. At the dense core, in contrast, we find a higher CH3CN column density (by a factor of $\sim 3$) compared to Gratier et al. (2013). This discrepancy is likely due to the choice of collisional rates: Gratier et al. (2013) adopted the rates from Green (1986), whereas Ben Khalifa et al. (2023) recently showed that those rates can be severely underestimated for some transitions, especially at low temperatures ($T_{\mathrm{kin}}=20$ K) such as those found in the dense core. Indeed, when we repeated our analysis using the older rates from Green (1986), our derived $N$ values matched those reported by Gratier et al. (2013). In the PDR, on the other hand, the choice of collisional coefficients has a less significant effect.
Concerning the shape of the $N$ profiles, we find that HNCO, CH3CN, CH2CO, and CH3CHO present higher column densities at the edge of the Horsehead nebula, with their peak near the PDR position. This suggests that their formation is enhanced in the presence of FUV radiation, consistent with the modern view of PDRs in which FUV photons can promote the production of certain molecules (e.g., Goicoechea et al. 2025). In contrast, we find that CH3OH and HC3N present higher column densities in deeper layers of the Horsehead nebula, with their peak near the dense core. These molecules are probably formed (or survive) more efficiently or readily desorbed into the gas-phase when shielded from FUV radiation. Lastly, H2CO stands out for its high abundance in both FUV-exposed and FUV-shielded layers of the Horsehead nebula, indicating that it can form efficiently under a range of physical conditions.
4 Discussion
4.1 Chemistry of O-bearing species
Figure 5 reveals clear differences in the column density distributions of molecular species with similar elemental compositions, likely reflecting distinct underlying chemical processes. In the case of O-bearing species, H2CO and CH3OH have their peak closer to the dense core position, whereas CH2CO and CH3CHO exhibit larger values near the PDR. This may seem surprising, considering that both molecular pairs are thought to be efficiently formed on dust grain surfaces through CO hydrogenation (see Herbst and van Dishoeck 2009, and references therein). However, the observed chemical differentiation may arise from the presence of an additional C atom that distinguishes the two groups of molecules.
Recent estimates of the binding energy of atomic C onto grain surfaces are reshaping the classical view proposed by early PDR models (e.g., Hollenbach et al. 2009). These models typically assumed a low binding energy ($E_{\mathrm{b}}\sim 800$ K, Hasegawa and Herbst 1993), implying that atomic C would freeze out onto grains only at very low temperatures ($T_{\mathrm{kin}}\lesssim 20$ K). However, more recent quantum chemical calculations suggest much higher binding energies ($E_{\mathrm{b}}>10^{4}$ K, Shimonishi et al. 2018; Duflot et al. 2021; Minissale et al. 2022), indicating that atomic C can deplete at significantly higher temperatures onto amorphous water ices, as those expected to coat grains within the interior of the Horsehead nebula (Schirmer et al. 2020; Abergel et al. 2024).
[…]
Near the PDR position, the presence of FUV radiation leads to the photodissociation of CO, resulting in the coexistence of CO, atomic C, and C+ at the illuminated edge of the Horsehead nebula (Philipp et al. 2006; Bally et al. 2018; Hernández-Vera et al. 2023). Atomic C can then be physisorbed onto H2O-rich ices on shallow binding sites, allowing its efficient diffusion for dust temperatures above $\sim$22 K (Tsuge et al. 2023), consistent with the conditions expected at the Horsehead PDR ($T_{\mathrm{dust}}\sim 30$ K, Goicoechea et al. 2009b). In contrast, in the dense core where dust temperatures are lower ($T_{\mathrm{dust}}\sim 20$ K, Goicoechea et al. 2009b), CO freeze-out becomes more relevant and atomic C diffusion is less efficient. Under these conditions, CO hydrogenation likely dominates, enhancing the formation of H2CO and CH3OH.
Interestingly, although H2CO is more abundant near the dense core, it also shows a significant contribution from the PDR, comparable to that of CH2CO and CH3CHO. This suggests that atomic C may also play a key role in the grain-surface formation of H2CO. In fact, an alternative to the traditional CO hydrogenation pathway involves the surface reaction C + H2O $\rightarrow$ H2CO, recently supported by both computational studies (Molpeceres et al. 2021) and laboratory experiments (Potapov et al. 2021). Consistently, Tsuge et al. (2023) found that a considerable fraction of the atomic C physisorbed onto H2O-rich ices reacts rapidly and forms H2CO.
An increased production of H2CO on grains enhances the cold formation pathways of CH2CO and CH3CHO, but also CH3OH (Potapov and Garrod 2024). However, under FUV irradiation, CH3OH tends to fragment into smaller radicals, which can subsequently lead to the formation of CH2CO and CH3CHO (Yocum et al. 2021), or even regenerate H2CO (Bertin et al. 2016). Furthermore, while CH3OH is predominantly formed via surface reactions on dust grains (Watanabe and Kouchi 2002; Fuchs et al. 2009), CH2CO, CH3CHO, and H2CO also follow viable gas-phase formation pathways (Ruiterkamp et al. 2007; Vazart et al. 2020; Ramal-Olmedo et al. 2021), as shown by recent detections of these species in diffuse gas lacking ice mantles (Gerin et al. 2025).
These pathways typically involve reactions between atomic O and hydrocarbon radicals—specifically, C2H3, C2H5, and CH3 for CH2CO, CH3CHO, and H2CO, respectively—highlighting another possible chemical link among the three species. Notably, atomic O has been detected at the edge of the Horsehead nebula (Goicoechea et al. 2009b), and the hydrocarbon radicals may originate from the photodissociation of closed-shell hydrocarbons like C2H6 and CH4 (Heays et al. 2017), which themselves are efficient products of grain-surface C chemistry (Kobayashi et al. 2017; Qasim et al. 2020; Tsuge et al. 2024). This would reflect a form of top-down hydrocarbon chemistry, complementing that previously proposed for PAHs at the edge of the Horsehead nebula (Pety et al. 2005; Guzmán et al. 2015).
Overall, whether O-bearing molecules form on dust grains or in the gas phase, the presence of atomic C under FUV-illuminated conditions appears to play a key role in driving their chemical complexity at relatively low temperatures ($T_{\mathrm{kin}}\lesssim 100$ K). In the dense core, by contrast, CO hydrogenation leading to H2CO and CH3OH becomes more important, since atomic C is less abundant and its diffusion on grain surfaces is less efficient. The UV-shielded conditions also limit the fragmentation of CH3OH, thereby influencing the formation of CH2CO and CH3CHO. Notably, the prominent detection of CH3OH—and O-bearing species in general—in the gas phase of the dense core may serve as direct evidence of chemical desorption, as previously suggested by Le Gal et al. (2017).
4.2 Chemistry of N-bearing species
Figure 5 also shows that N-bearing species display notably different column density distributions. While CH3CN shows a steep profile peaking near the PDR position, HC3N presents a more gradual increase toward the dense core. In contrast, HNCO exhibits similarities with the O-bearing molecules, as previously indicated by its integrated intensity distribution (see Fig. 3). In line with our interpretation for the O-bearing species, we propose that the chemistry of N-bearing molecules in FUV-irradiated regions is likewise influenced by atomic C chemistry.
According to the literature, CH3CN can be initiated in the gas phase through the radiative association reaction $_{3}^{+} + HC$N $\rightarrow$ CH3CNH+ + $h\nu$, followed by dissociative recombination with electrons (Giani et al. 2023). Nevertheless, its enhanced column density near the PDR may indicate a contribution from dust grain surface chemistry, as previously suggested by Gratier et al. (2013) and supported by experimental evidence showing that CH3CN undergoes more efficient photodesorption than molecules such as CH3OH (Basalgète et al. 2021). One proposed grain-surface formation pathway involves the reaction between CH3 and CN radicals to form CH3CN (Garrod et al. 2008; Enrique-Romero and Lamberts 2025), with both radicals potentially originating from the diffusion of atomic C or the photodissociation of larger species.
Another possible route of CH3CN on grains is through the hydrogenation of C2N (Walsh et al. 2014). This molecule is thought to form primarily in the gas-phase via the reaction C2H + N $\rightarrow$ H + C2N, followed by freeze-out onto dust grains (Loomis et al. 2018). However, there are grain-surface reactions like C + CN or C2 + N $\rightarrow$ C2N listed as potential pathways in the Kinetic Database for Astrochemistry (KIDA; Wakelam et al. 2012, 2024) that have yet to be experimentally confirmed. Additionally, Canta et al. (2023) probed that the formation of CH3CN is also possible due to the photoprocessing of ice mixtures composed by NH3 and hydrocarbons of the form C2Hx, where the latter are possibly produced by atomic C on dust grains.
[…]
In the case of HC3N, it is generally accepted that this molecule lacks efficient grain-surface formation pathways (Bergner et al. 2017, 2018). This may account for the absence of a filamentary emission structure at the PDR, in contrast to most of the species discussed in this work, which likely receive significant contributions from dust grain chemistry. In the gas phase, the formation of HC3N is primarily driven by hydrocarbon reactions such as C2H2 + CN and C2H + HCN $\rightarrow$ HC3N + H (Fukuzawa and Osamura 1997; Taniguchi et al. 2024). A less dominant, yet non-negligible, route involves C3H2 + N $\rightarrow$ HC3N (Murillo et al. 2022). In the Horsehead nebula, these hydrocarbon precursors (e.g., C2H, C2H2, and C3H2) are believed to originate mainly from gas-phase reactions initiated by C+ and CH, along with PAH photochemistry (Guzmán et al. 2015). However, an alternative origin linked to products of grain-surface C chemistry, such as CH4 (Lamberts et al. 2022; Tsuge et al. 2024), may also contribute (e.g., Hassel et al. 2008; Sakai and Yamamoto 2013; Loison et al. 2017).
PdBI maps of C2H and $c$-C3H2 provide additional evidence that these species are chemically linked to HC3N. Guzmán et al. (2015) noted that the emission of these hydrocarbons follows the same two-filament structure recently revealed by high angular resolution observations of HCO+, located in the vicinity of the PDR and dense core (Hernández-Vera et al. 2023). At the PDR, HC3N does not display a noticeable filament, likely because it is more efficiently destroyed than formed—consistent with its rapid photodissociation relative to other N-bearing species such as CH3CN (Le Gal et al. 2019), as well as the lack of efficient grain-surface formation pathways. In contrast, within the filament near the dense core, all three molecules trace the same emitting region, although C2H appears to be depleted where $c$-C3H2 and HC3N are brighter. Moreover, at the exact position of the dense core—where HCO+ and DCO+ reach their emission peaks (Pety et al. 2007; Hernández-Vera et al. 2023)—HC3N shows a slight depletion, in line with its predicted rapid destruction via ion–molecule reactions (Taniguchi et al. 2024).
Lastly, although HNCO is composed by nitrogen and oxygen, we decided to consider it as a N-bearing molecule due to the key role of atomic N during its formation under FUV-irradiated conditions. Despite its relative simplicity, the formation pathways of HNCO remain poorly constrained. Current models suggest a combined contribution from both gas-phase and grain-surface chemistry, though disentangling their respective roles is challenging (Tideswell et al. 2010; Quan et al. 2010). In both cases, the radical NCO is considered a key intermediate (Marcelino et al. 2010). In the gas phase, HNCO can be formed via protonation of NCO by proton donors like $_{3}^{+}, H$3O+, and HCO+, followed by further ion–molecule reactions and eventual dissociative recombination (Velilla Prieto et al. 2015; Cernicharo et al. 2024). On grain surfaces, hydrogenation of NCO has been proposed as an efficient alternative as well (Quan et al. 2010; Belloche et al. 2017).
However, the chemistry of NCO itself is also intricate. Its formation is typically modeled in the gas phase through reactions such as N + HCO → NCO + H (Quan et al. 2010; Cernicharo et al. 2024), CN + O2 → NCO + O (Marcelino et al. 2018; Cernicharo et al. 2024), and CN + OH → NCO + H (Velilla Prieto et al. 2015). Once produced, NCO may accrete onto dust grains (Quan et al. 2010), although direct formation on icy mantles is also possible via reactions between CO and atomic N (Fedoseev et al. 2015; Mignon et al. 2017), or through pathways involving CO photodissociation products such as O + CN and C + NO (Belloche et al. 2017). Notably, the latter reaction may benefit from efficient surface diffusion of atomic C.
Figure 6: Schematic summary of the proposed scenario at the edge of the Horsehead nebula. The two crosses indicate the PDR and dense core positions, as in Fig. 2. At the PDR, the moderate FUV field enables the coexistence of C and CO, together with a sufficiently high $T_{\mathrm{dust}}$ to allow efficient diffusion of atomic C on grain surfaces. These conditions may favor the formation of H2CO, CH2CO, CH3CHO, HNCO, and CH3CN, likely through a combination of gas-phase and grain-surface pathways. Interestingly, for the two species with the lowest PDR abundances, CH3OH and HC3N, previous studies suggest ineffective photodesorption and inefficient grain-surface formation, respectively. In the case of the dense core, the FUV-shielded conditions limit CO photodissociation, and the low $T_{\mathrm{dust}}$ makes the diffusion of atomic C on grains inefficient. These conditions may favor the grain-surface formation of H2CO and CH3OH, and also contribute to the prevalence of HC3N in the gas phase, as it is less efficiently photodissociated.
Another important formation route for HNCO involves the simultaneous hydrogenation and UV processing of gas-phase NO after its accretion onto CO, H2CO, and CH3OH-rich ices (Fedoseev et al. 2016). This mechanism is particularly relevant given the abundant detection of NO at the edge of the Horsehead nebula (WHISPER collaboration, private communication). Laboratory experiments also suggest that the irradiation of ice mixtures containing other N-bearing species (e.g., N2, NH3, and HCN) may contribute to HNCO formation through complementary pathways (Jiménez-Escobar et al. 2014; Martín-Doménech et al. 2020, 2024). Such processes may help explain the elevated column densities observed near the PDR position. On the other hand, hydrogenation products of atomic N, such as NH and NH2, can efficiently react with CO on grain surfaces to form HNCO (Fedoseev et al. 2015; Belloche et al. 2017), in a manner analogous to the surface chemistry proposed for O-bearing species. This could account for the relatively high column densities of HNCO near the dense core, compared to other N-bearing molecules.
Similar to O-bearing species, these findings suggest that the chemistry of N-bearing molecules is closely linked to that of atomic C under FUV-irradiated conditions, both in the gas phase and on dust grain surfaces. Nonetheless, atomic N and its derivative species may also play a significant role, particularly in the case of HNCO, which contains fewer C atoms in its structure. To achieve a more comprehensive understanding of the formation and evolution of HNCO and N-bearing molecules in general, further dedicated studies targeting smaller N-bearing species (e.g., HCN, HNC, HNO, NCO, CN, and NO) in combination with detailed chemical modeling, are likely necessary.
Figure 6 presents a schematic overview of the physicochemical processes proposed in this work. Despite the experimental evidence and observational signposts that atomic C may contribute to the formation of COMs, modeling its surface chemistry remains challenging. A recent study by Lis et al. (2025), using the Meudon PDR code (Le Petit et al. 2006) wrapped in a locally circular 2D geometry, found difficulties in reproducing the observed water vapor emission in the Horsehead nebula, suggesting that current models may be missing key grain-surface or photochemical processes. This highlights the need for further refinement of surface chemistry treatments, which appear to be fundamental for accurately modeling the abundances of COMs.
4.3 Implications in other UV-irradiated environments
The insights gained from the Horsehead nebula can inform our understanding of the formation of COMs in other UV-irradiated environments. Just as the Horsehead PDR serves as a benchmark for regions where molecular gas is exposed to moderate FUV radiation, the Orion Bar PDR has traditionally been regarded as a prototypical case for strongly FUV-irradiated molecular gas (Tielens et al. 1993; Hogerheijde et al. 1995; Goicoechea et al. 2016).
As a consequence of the higher FUV flux in the Orion Bar PDR ($G_{0}>10^{4}$, Marconi et al. 1998; Peeters et al. 2024), the degree of photodissociation is expected to be significantly greater than in the Horsehead PDR. Furthermore, dust grains in the Orion Bar are warmer ($T_{\mathrm{dust}}\simeq 40$$80$ K, Arab et al. 2012), which inhibits the formation of substantial ice mantles and thereby reduces the efficiency of grain surface chemistry. Indeed, we have found that most of the molecules analyzed in this work are less abundant in the gas phase of the Orion Bar PDR (Cuadrado et al. 2017) compared to the Horsehead PDR, with the remarkable exception of H2CO and CH3OH (see Table 4).
Given that CH3OH is predominantly formed on grains—with H2CO as one of its main precursors—a non-negligible grain-surface contribution deep inside the Orion Bar cannot be ruled out. Cuadrado et al. (2017) showed that pure gas-phase models fail to reproduce the observed abundances of H2CO and CH3OH, suggesting that the formation of organics on warm grains with limited ice mantles may still be feasible. Nevertheless, given the high abundance of C at the edge of the Orion Bar (Goicoechea et al. 2025) and the elevated dust temperatures, an efficient surface diffusion of atomic C on grains should play a significant role, but apparently is not the case. Hence, in contrast to the Horsehead PDR, where C can be sufficiently abundant on grains to compete with CO hydrogenation, the Orion Bar PDR appears to process ices dominated by CO rather than C, resembling those expected in the Horsehead dense core. This is, however, unlikely to occur in situ in the Bar owing to the high temperatures—though it may operate in dense cloud interiors, as evidenced by spatially resolved CH3OH observations (Leurini et al. 2010).
A possible explanation to these contrasting behaviors in surface chemistry products may lie in PDR dynamics. Cuadrado et al. (2017) already speculated that, in the Orion Bar, the advection of material from the cold, UV-shielded interior of the molecular cloud to the warmer PDR edge is plausible in the presence of considerable dynamical effects. Supporting this, Goicoechea et al. (2016) demonstrated that dynamical processes can significantly shape the cloud structure in the Orion Bar. In the case of the Horsehead, although dynamical effects are expected to be even stronger than in high-excitation PDRs (Maillard et al. 2021), the present results, together with the absence of photoablative molecular gas flows and the lower compression factors (Hernández-Vera et al. 2023), could suggest a less prominent dynamical influence compared to the Orion Bar, at least concerning COMs formation. Nevertheless, higher angular resolution observations are required to confidently assess this scenario.
[…]
Our spatially resolved analysis of complex organics in the Horsehead nebula can also have important implications for protoplanetary disks. Driven primarily by UV radiation and X-rays from the central star, the chemical composition observed in the cold outer regions of nearby disks qualitatively resembles that expected for classical PDRs (Öberg and Bergin 2021; Öberg et al. 2023). Rich in nitriles and hydrocarbons, yet notably deficient in large O-bearing organics, these disk regions are typically characterized by a high C/O ratio, a scenario similarly proposed for the Horsehead PDR. Furthermore, recent ALMA observations have uncovered extended [C I] emission in protoplanetary disks (Booth et al. 2023; Law et al. 2023; Urbina et al. 2024), possibly indicating that atomic C could have a relevant yet largely unexplored chemical role within these environments.
In this work, we find that CH3OH is not the dominant O-bearing species in the gas phase under moderate FUV irradiation, which may help explain its elusive detection in cold Class II disks (e.g., Carney et al. 2019; Loomis et al. 2020; Yamato et al. 2024). In contrast, H2CO is abundantly detected in disks (e.g., Pegues et al. 2020; Guzmán et al. 2021), yet its potential formation via grain-surface reactions involving atomic C remains unexplored. If efficient, this pathway would increase CH3OH production, whose photofragmentation might drive further chemical complexity, as observed in the Horsehead PDR. Similarly, N-bearing species like CH3CN and HC3N are also common in disks (Bergner et al. 2018; Ilee et al. 2021), with their formation likely tied to UV-driven chemistry (Calahan et al. 2023).
Although dust grains in protoplanetary disks typically settle toward the cold and UV-shielded midplane (Blum and Wurm 2008), turbulent mixing can transport icy material into more irradiated layers (Öberg et al. 2015; Bergner and Ciesla 2021; Ligterink et al. 2024). Models including this process predict notable effects for species like CH2CO, CH3CHO, and HNCO (Semenov and Wiebe 2011), which—despite their prebiotic relevance—are rarely targeted in disk studies. Recent observations by Yamato et al. (2024) also reveal abundant CH3OCH3 in the cold disk MWC 480 due to in situ ice processing. Altogether, these results suggest that PDR-like chemistry provides valuable context for disk studies, motivating the search for tracers beyond CH3OH.
5 Summary and Conclusions
We have investigated the spatial distribution of simple and complex organic molecules at the FUV-irradiated edge of the Horsehead nebula. Using ALMA-7m interferometry combined with IRAM 30m single-dish data, we have mapped species such as C17O, HNCO, HC3N, CH3CN, CH2CO, and CH3CHO for the first time in this source, achieving an angular resolution of approximately $\sim$$15^{\prime\prime}$. Our analysis also incorporates deep single-dish observations of C17O and C18O from the WHISPER survey, along with PdBI maps of H2CO and CH3OH from Guzmán et al. (2013). Integrated line intensity profiles were extracted along the direction of the FUV radiation, and radiative transfer modeling was used to constrain the gas physical conditions and molecular column densities. The main results and conclusions are summarized as follows:
  1. 1.
    The column density of C17O in the Horsehead nebula has been constrained for the first time. Using previous estimates of $N(\mathrm{H_{2}})$, an empirical C17O-to-H2 conversion factor of $X_{\mathrm{C^{17}O}}=N(\mathrm{H_{2}})/N(\mathrm{C^{17}O})\sim(5\pm 1)\times 10^{6}$ is obtained. The newly derived ratio $N(\mathrm{C^{18}O})/N(\mathrm{C^{17}O})\sim 3.5$ further confirms that this $X_{\mathrm{C^{17}O}}$ value is consistent with previous estimates from other CO isotopologues.
  2. 2.
    Application of the $X_{\mathrm{C^{17}O}}$ factor to the $N(\mathrm{C^{17}O})$ profiles yields empirical $n_{\mathrm{H_{2}}}$ profiles, corroborating the presence of a steep density gradient at the edge of the Horsehead nebula, in agreement with prior models and observations.
  3. 3.
    Simultaneous fitting of multiple $K$-ladder transitions of CH3CN enabled the determination of empirical $T_{\mathrm{kin}}$ profiles, confirming previous temperature estimates at both the dense core and PDR positions obtained from various independent tracers.
  4. 4.
    From the $n_{\mathrm{H_{2}}}$ and $T_{\mathrm{kin}}$ profiles, we derived an empirical profile of $P_{\mathrm{th}}$. The high thermal pressure values, together with their relatively small variations across the distance from the FUV radiation source, confirm the compressed and isobaric nature of the Horsehead PDR.
  5. 5.
    Except for HC3N and CH3OH, most of the analyzed molecules—specifically H2CO, CH2CO, CH3CHO, HNCO, and CH3CN—show a clear increase in column density toward the PDR, suggesting that exposure to FUV irradiation favors their gas-phase abundance, contrary to classical PDR expectations. In the Horsehead, this likely points to a significant contribution from grain surface chemistry, given that HC3N is not efficiently formed on grains, while CH3OH, although formed on grains, is not efficiently desorbed.
  6. 6.
    Given that atomic C is expected to be abundant in PDRs due to CO photodissociation, and considering the widespread detection of O- and N-bearing COMs, the Horsehead nebula serves as a compelling example of how surface diffusion of atomic C on grains could enhance chemical complexity in the ISM, as recently suggested by laboratory experiments. Atomic C may contribute either by directly forming COMs on grain surfaces or by producing precursors that are subsequently desorbed, driving further chemical evolution in the gas phase.
Overall, compared to regions exposed to stronger FUV radiation fields, such as the Orion Bar, the Horsehead PDR appears to offer an optimal combination of temperature and UV intensity to promote the efficient formation of COMs through a complex interplay of gas-phase and grain-surface chemistry. Disentangling the relative contributions of these pathways will require higher angular resolution observations and/or James Webb Space Telescope (JWST) ice studies toward stars located behind the Horsehead. Our results further suggest that similar chemical processes may operate in other cold, UV-irradiated environments with abundant CO and H2O ices, such as the molecular layers of protoplanetary disks. However, additional chemical modeling and new observations will be essential to assess the broader applicability of these findings to other sources.
Acknowledgements.
The authors thank the referee for the constructive comments that improved the content of this work. This paper makes use of the following ALMA data: ADS/JAO.ALMA#2016.2.00027.S. ALMA is a partnership of ESO (representing its member states), NSF (USA) and NINS (Japan), together with NRC (Canada), NSTC and ASIAA (Taiwan), and KASI (Republic of Korea), in cooperation with the Republic of Chile. The Joint ALMA Observatory is operated by ESO, AUI/NRAO and NAOJ. The National Radio Astronomy Observatory is a facility of the National Science Foundation operated under cooperative agreement by Associated Universities, Inc. Data reduction was carried out at the IRAM ARC node. C.H.-V. acknowledges support from the National Agency for Research and Development (ANID) – Scholarship Program through the Doctorado Nacional grant no. 2021-21212409. C.H.-V. and V.V.G. acknowledge support from the ANID – Millennium Science Initiative Program – Center Code NCN2024_001. V.V.G also gratefully acknowledges support from FONDECYT Regular 1221352, and ANID CATA-BASAL project FB210003. J.P. and M.G. acknowledge support by the French Agence Nationale de la Recherche through the DAOISM grant ANR-21-CE31-0010 and by the Thematic Action “Physique et Chimie du Milieu Interstellaire” (PCMI) of INSU Programme National “Astro”, with contributions from CNRS Physique & CNRS Chimie, CEA, and CNES. K.T.W. acknowledges support from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (Grant agreement no. 883867, project EXWINGS). J.R.G. thanks the Spanish MCINN for funding support under grant PID2023-146667NB-I00.
Appendix A Observational details
Table 2 summarizes the observational details of the spectral setup used in the ALMA-7m observations.
Table 2: Spectral setup of the ALMA-7m observations [^10] Columns from left to right: baseband number, rest frequency (GHz) at the centre of the spectral window, targeted species, channel spacing (kHz), bandwidth (MHz), and the central frequency (GHz) of the TDM spectral window used in calibration.
$B$ $\nu_{\rm rest}$ Species $\Delta\nu$ BW $\nu_{\rm TDM}$
(GHz) (kHz) (MHz) (GHz)
1 98.00 $l$-C3H 30.5 62.5 98.32
98.86 CH3CHO 30.5 62.5
2 100.09 CH2CO, HC3N 30.5 62.5 100.28
100.53 CH3NC 30.5 62.5
3 109.91 HNCO 30.5 62.5 110.32
110.38 CH3CN 30.5 62.5
4 112.25 CH3CHO 244 250 112.32
112.36 C17O 244 250
112.70 Continuum 244 500
Appendix B Bayesian fitting results
B.1 WHISPER CO isotopologues observations
Figure 7 presents the best-fit models accounting for the hyperfine structure of the C17O (top panels) and C18O (bottom panels) J=10J=1-0 and 212-1 rotational transitions, based on the WHISPER data analysis described in Sect. 3.2.1. The best-fit parameters and their statistical uncertainties for the PDR and dense core positions are shown in the middle of each panel. Averaging the results for the PDR and dense core yields $T_{\mathrm{ex}}(\mathrm{C^{17}O})\sim 15$ K and $N(\mathrm{C^{18}O})/N(\mathrm{C^{17}O})\sim 3.6$, which are the values adopted in Sect. 3.2.1 for the posterior analysis of the ALMA-7m+30m combined dataset.
Figure 7: C17O (top panels) and C18O (bottom panels) spectra at the PDR position ($\delta x\approx 15\hbox{$.\!\!^{\prime\prime}$}1,\>\delta y\approx 1\hbox{$.\!\!^{\prime\prime}$}2$) and dense core position ($\delta x\approx 44\hbox{$.\!\!^{\prime\prime}$}7,\>\delta y\approx 16\hbox{$.\!\!^{\prime\prime}$}5$), as observed in the Horsehead WHISPER survey. The best-fit models are overlaid in red. The best-fit excitation conditions and associated uncertainties, shown in the middle of each panel, correspond to the 16th, 50th, and 84th percentiles of the posterior distributions.
B.2 ALMA-7m+30m C17O observations
Figure 8 presents zeroth-moment maps of the hyperfine components of the C17O J=10J=1-0 rotational transition from the combined ALMA-7m+30m dataset. Only two maps are displayed, as the two hyperfine components at $112.359$ GHz are blended (see Table 3). The green and cyan squares mark the PDR and dense core positions, respectively, where the spectra were extracted for the calibration of the C17O-to-H2 conversion factor.
Below the maps, the corresponding spectra and best-fit models are displayed, with the derived parameters and their associated uncertainties indicated, following a similar format as in Fig. 7. By assuming $T_{\mathrm{ex}}(\mathrm{C^{17}O})=15$ K, we derive column densities of $N(\mathrm{C^{17}O})=(4.7\pm 0.2)\times 10^{15}$ cm-2 for the PDR and $N(\mathrm{C^{17}O})=(5.2\pm 0.2)\times 10^{15}$ cm-2 for the dense core. These values are slightly higher than those obtained with WHISPER, likely due to beam dilution. As a sensitivity test, we repeated the fitting procedure using the $T_{\mathrm{ex}}$ values estimated from WHISPER at both positions (see Fig. 7), finding no substantial changes in the derived $N(\mathrm{C^{17}O})$ or FWHM values.
These results, together with $N(\mathrm{H_{2}})$ estimates from Guzmán et al. (2014), were used to calibrate the C17O-to-H2 conversion factor discussed in Sect. 3.2.1. Equivalent plots were generated (but are not shown) for the remaining $\delta x$ positions at the heights of the PDR ($\delta y\sim 1\hbox{$.\!\!^{\prime\prime}$}2$) and dense core ($\delta y\sim 16\hbox{$.\!\!^{\prime\prime}$}5$), enabling the computation of the $n_{\mathrm{H_{2}}}$ profiles as described in Sect. 3.2.2.
Figure 8: Top row: ALMA-7m+30m zeroth-moment maps of the hyperfine components of the C17O J=10J=1-0 rotational transition resolved at the Horsehead edge. The rotation angle and red vertical line are the same as in Fig. 2. Contours are $5,10,20,30,50,100]\times\sigma$, where $\sigma$ is the zeroth-moment rms listed in Table 1. The dense core and PDR positions are marked with cyan and green squares, respectively, indicating the regions where the spectra were extracted. The beam size and a scale bar indicating $4000$ au are shown in the bottom left and bottom right corner, respectively, of each panel. Bottom row: C17O spectra extracted at the dense core and PDR positions, shown using matching colors to the map markers. The best-fit models are overlaid in red. The best-fit excitation conditions and associated uncertainties, presented in the top left corner of each panel, correspond to the 16th, 50th, and 84th percentiles of the posterior distributions.
B.3 ALMA-7m+30m CH3CN observations
Figure 9 presents the best-fit models to the integrated intensity profiles of the various CH3CN J=65J=6-5 ($K=0,1,$ and $2$) transitions, extracted at the vertical positions of the dense core (left panel) and the PDR (right panel), following the procedure described in Sect. 3.2.3. Near the PDR ($\delta x<30^{\prime\prime}$), the $K=0,1,$ and $2$ transitions were jointly fit. However, toward the dense core ($\delta x\geq 30^{\prime\prime}$), the integrated intensity of the $K=2$ transition was used only as an upper limit due to observed line profile asymmetries and broader line widths relative to the other $K$-ladder transitions. As the $K=2$ line corresponds to the highest $E_{u}$ among the analyzed CH3CN transitions (see Table 3), we speculate that its emission does not originate in the cold dense core, but rather arises from warmer gas along the line of sight, a phenomenon previously reported in the Horsehead for tracers such as CS and HCO (Goicoechea et al. 2006; Gerin et al. 2009). It is also plausible that this warmer, more diffuse component has different kinematics compared to the quiescent dense core, which could explain not only the line broadening but also the asymmetric deviations from a purely Gaussian line profile.
Figure 9: Integrated intensity profiles of the CH3CN J=65J=6-5 ($K=0,1,$ and $2$) transitions at the $\delta y$ positions of the dense core (left panel) and the PDR (right panel), as obtained from the ALMA-7m+30m observations. Each color corresponds to a different $K$-ladder transition, as indicated by the labels in the right panel. Observed data points are shown as circles, and the best-fit models are overlaid as solid lines in the corresponding colors. The red vertical line, the dotted vertical lines, and the horizontal gray bar are the same as in the left panels of Fig. 3.
Appendix C Molecular data
Table 3 summarizes the spectroscopic constants and the critical densities for all the molecular lines analyzed throughout this work.
Table 3: Molecular data of the transitions analyzed in this work. [^11] This covers all the lines analyzed throughout this work, including the legacy data described in Sect. 2.4. All molecular data are taken from the Cologne Database for Molecular Spectroscopy (CDMS; Endres et al. 2016) except for CH3CHO, which is from the Jet Propulsion Laboratory database (JPL; Pickett et al. 1998).
Species Transition Frequency $E_{u}$ $\log_{10}A_{ul}$ $g_{u}$ $n_{\mathrm{crit}}$aaaaaaFor those species with available collisional rates in the Leiden Atomic and Molecular Database (LAMDA; Schöier et al. 2005), the critical density ($n_{\mathrm{crit}}$) is estimated as the density at which the rate of spontaneous radiative de-excitation ($A_{ul}$) equals the rate of collisional de-excitation ($\gamma_{ul}$) for temperatures between 206020-60 K.
(GHz) (K) (s-1) (cm-3)
C17O J=10J=1-0, F=3/25/2F=3/2-5/2 112.359 5.39 -7.174 4 $(2.0-2.1)\times 10^{3}$
J=10J=1-0, F=7/25/2F=7/2-5/2 112.359 5.39 -7.174 8 $(2.0-2.1)\times 10^{3}$
J=10J=1-0, F=5/25/2F=5/2-5/2 112.360 5.39 -7.174 6 $(2.0-2.1)\times 10^{3}$
J=21J=2-1, F=3/25/2F=3/2-5/2 224.713 16.18 -6.715 4 $(1.0-1.1)\times 10^{4}$
J=21J=2-1, F=5/25/2F=5/2-5/2 224.714 16.18 -6.407 6 $(1.0-1.1)\times 10^{4}$
J=21J=2-1, F=7/25/2F=7/2-5/2 224.714 16.18 -6.384 8 $(1.0-1.1)\times 10^{4}$
J=21J=2-1, F=9/27/2F=9/2-7/2 224.714 16.18 -6.192 10 $(1.0-1.1)\times 10^{4}$
J=21J=2-1, F=1/23/2F=1/2-3/2 224.714 16.18 -6.192 2 $(1.0-1.1)\times 10^{4}$
J=21J=2-1, F=3/23/2F=3/2-3/2 224.715 16.18 -6.347 4 $(1.0-1.1)\times 10^{4}$
J=21J=2-1, F=5/27/2F=5/2-7/2 224.715 16.18 -7.292 6 $(1.0-1.1)\times 10^{4}$
J=21J=2-1, F=7/27/2F=7/2-7/2 224.715 16.18 -6.639 8 $(1.0-1.1)\times 10^{4}$
J=21J=2-1, F=5/23/2F=5/2-3/2 224.715 16.18 -6.699 6 $(1.0-1.1)\times 10^{4}$
C18O J=10J=1-0 109.782 5.27 -7.203 3 $(1.8-1.9)\times 10^{3}$
J=21J=2-1 219.560 15.81 -6.221 5 $(0.9-1.0)\times 10^{4}$
H2CO $J_{K_{a},K_{c}}=2_{02}-1_{01}$ 145.603 10.48 -4.107 5 $(0.9-1.1)\times 10^{6}$
CH3OH $J_{K}=3_{-1}-2_{-1}$ (E) 145.097 19.51 -4.960 7 $(8.4-9.1)\times 10^{4}$
CH2CObbbbbbAlthough CH2CO and CH3CHO lack available collisional rates, their emission is likely sub-thermal ($n_{\mathrm{H}_{2}}<n_{\mathrm{crit}}$), as suggested by the low $T_{\mathrm{ex}}$ values reported by Guzmán et al. (2014), which are adopted in this work. $J_{K_{a},K_{c}}=5_{15}-4_{14}$ 100.095 27.46 -4.988 33 -
CH3CHObbbbbbAlthough CH2CO and CH3CHO lack available collisional rates, their emission is likely sub-thermal ($n_{\mathrm{H}_{2}}<n_{\mathrm{crit}}$), as suggested by the low $T_{\mathrm{ex}}$ values reported by Guzmán et al. (2014), which are adopted in this work. $J_{K_{a},K_{c}}=5_{14}-4_{13}$ (E) 98.863 16.59 -4.508 22 -
$J_{K_{a},K_{c}}=6_{16}-5_{15}$ (E) 112.255 21.21 -4.330 26 -
$J_{K_{a},K_{c}}=6_{16}-5_{15}$ (A) 112.249 21.13 -4.331 26 -
HNCO $J_{K_{a},K_{c}}=5_{05}-4_{04}$ 109.906 15.82 -4.756 11 $(2.4-2.5)\times 10^{5}$
HC3N J=1110J=11-10 100.076 28.82 -4.110 23 $(1.5-1.7)\times 10^{6}$
CH3CN $J_{K}=6_{0}-5_{0}$ (A) 110.384 18.54 -3.954 26 $(1.9-2.4)\times 10^{6}$
$J_{K}=6_{1}-5_{1}$ (E) 110.381 25.69 -3.966 26 $(2.0-2.5)\times 10^{6}$
$J_{K}=6_{2}-5_{2}$ (E) 110.375 47.12 -4.005 26 $(1.6-2.0)\times 10^{6}$
Appendix D Abundances at the PDR and dense core positions
Figure 10 presents the column density profiles derived at the vertical position of the dense core, analogous to those shown for the vertical position of the PDR in Sect. 3.2.5. For comparison with the Orion Bar PDR, we extracted the column density values specifically at the $\delta x$ positions of the dense core ($\delta x\approx 44\hbox{$.\!\!^{\prime\prime}$}7$, from Fig. 10) and the PDR ($\delta x\approx 15\hbox{$.\!\!^{\prime\prime}$}1$, from Fig. 5) to estimate their abundances with respect to total hydrogen nuclei. The resulting values, together with those for the Orion Bar reported in the literature, are listed in Table 4.
Figure 10: Same as Fig. 5, but at the $\delta y$ position of the dense core.
Table 4: Abundances with respect to total hydrogen nuclei ($N(\mathrm{X})/N_{\mathrm{H}}$, in units of $10^{-10}$). [^12] The total hydrogen nuclei column density is estimated as $N_{\mathrm{H}}=N(\mathrm{X})/(N(\mathrm{H})+2N(\mathrm{H}_{2}))$.
Species HorseheadaaaaaaThis work. Abundances calculated using $N_{\mathrm{H}}=3.8\times 10^{22}$ cm-2 (PDR) and $N_{\mathrm{H}}=6.4\times 10^{22}$ cm-2 (dense core). Orion BarbbbbbbValues from Cuadrado et al. (2017). Abundances calculated using $N_{\mathrm{H}}=6.3\times 10^{22}$ cm-2.
PDR Dense core PDR
H2CO $3.4^{+0.5}_{-0.5}$ $3.0^{+0.4}_{-0.3}$ $9.0$
CH3OH $0.4^{+0.7}_{-0.3}$ $1.8^{+0.3}_{-0.3}$ $5.0$
CH2CO $4.7^{+0.3}_{-0.3}$ $1.4^{+0.1}_{-0.1}$ $0.9$
CH3CHO $3.9^{+0.2}_{-0.1}$ $0.9^{+0.1}_{-0.1}$ $0.8$
CH3CN $1.3^{+0.1}_{-0.1}$ $^{+0.1}_{-0.1} 0.20.2$
HC3N $0.3^{+0.1}_{-0.1}$ $0.3^{+0.1}_{-0.1}$ $0.1$
HNCO $2.5^{+0.2}_{-0.1}$ $0.8^{+0.1}_{-0.1}$ $0.2$